# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v2

@@ 8 Summary
We offer a formal treatment of asymptotic and compound e-values by giving them rigorous definitions. We provide several prototypical constructions, and explain why compound e-values are a fundamental concept in multiple testing. We demonstrate several ways of constructing approximate or exact compound e-values, for example,
@@
    by leveraging any FDR controlling procedure,
@@
    through the compound likelihood ratio (a ratio of mixture likelihoods) which yields log-optimal simple separable compound e-values,
@@
    through compound universal inference,
@@
    or by calibration of average significance controlling p-values.
@@
We briefly point out that the term compound p-values has previously been used in a way that is distinct from Definition 4.1. Habiger and Peña (2014) considered a setting akin to the sequence model in Section 7 wherein $X=(X_{k}:k\in\mathcal{K})$ and in principle $X_{k}$ can be used to test $H_{k}$ for all $k\in\mathcal{K}$. Let $P_{1},\dots,P_{K}$ be $[0,1]$-valued random variables. Habiger and Peña (2014) called these “compound p-values” if they are bona fide p-values (that is $\mathbb{P}(P_{k}\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}_{k}$) and depend also on data for other hypotheses $(X_{\ell}:\ell\neq k)$, i.e., $P_{k}=P_{k}(X)$ rather than $P_{k}=P_{k}(X_{k})$. In other words, “compound p-values” are not separable as per Definition 7.1, and so could be called “non-separable p-values” instead. The definition treats separable p-values as the default data analysis choice (which is reasonable given how p-values are often used in practice), and one needs a qualifier to refer to non-separable p-values. By contrast, we prefer to think of (compound) p-values and e-values as being non-separable by default (given their flexibility and generality), and to use an explicit qualifier for the separable case. Moreover, we use the term compound in relation to the averaging (compounding) operation inherent in Definition 4.1.

### v3

@@ 8 Summary
We offer a formal treatment of asymptotic and compound e-values by giving them rigorous definitions. We provide several prototypical constructions, and explain why compound e-values are a fundamental concept in multiple testing. We demonstrate several ways of constructing bona fide, approximate, or asymptotic compound e-values, for example,
@@
    by leveraging any FDR controlling procedure;
@@
    through the optimal discovery compound e-values (which are ratios of mixture likelihoods) and empirical Bayes approximations thereof;
@@
    through localized compound universal inference;
@@
    or by calibration of (approximate) compound p-values.
@@
We briefly point out that the term compound p-values has previously been used in a way that is distinct from Definition 2.1. Habiger and Peña (2014) considered a setting akin to the sequence model in Section 5 wherein $X=(X_{k}:k\in\mathcal{K})$ and in principle $X_{k}$ can be used to test $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$. Let $P_{1},\dots,P_{K}$ be $[0,1]$-valued random variables. Habiger and Peña (2014) called these “compound p-values” if they are bona fide p-values (that is $\mathbb{P}(P_{k}\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}_{k}$) and depend also on data for other hypotheses $(X_{\ell}:\ell\neq k)$, i.e., $P_{k}=P_{k}(X)$ rather than $P_{k}=P_{k}(X_{k})$. In other words, “compound p-values” are not separable as per Definition 5.1, and so could be called “non-separable p-values” instead. The definition treats separable p-values as the default data analysis choice (which is reasonable given how p-values are often used in practice), and one needs a qualifier to refer to non-separable p-values. By contrast, we prefer to think of (compound) p-values and e-values as being non-separable by default (given their flexibility and generality), and to use an explicit qualifier for the separable case. Moreover, we use the term compound in relation to the averaging (compounding) operation inherent in Definition 2.1.
Reproducibility.
We provide code to reproduce the numerical results of this paper on Github:
https://github.com/nignatiadis/compound-evalues-paper
Acknowledgments.
This work was completed in part with resources provided by the University of Chicago’s Research Computing Center.